OpenStudy (anonymous):

solve: 6-4(2x+1) =5(3-2x)

OpenStudy (anonymous):

\[6-4(2x+1) =5(3-2x)\] distribute first and start with \[6-8x-4=15-10x\] then combine like terms on the left

OpenStudy (anonymous):

let me know what you get it will bet three steps after that

OpenStudy (anonymous):

is the answer 2x-13?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh i see \[-8x+2=15-10x\] add \(10x\) to both sides, get \[2x+2=15\]

OpenStudy (anonymous):

the subtract \(2\) and get \[2x=13\]

OpenStudy (anonymous):

finally divide by \(2\) and finish with \[x=\frac{13}{2}\]

OpenStudy (anonymous):

Thanks!